### v1

# Learning General Representation of 12-Lead Electrocardiogram With a Joint-Embedding Predictive Architecture

Abstract
We propose a self-supervised learning method for 12-lead Electrocardiogram (ECG) analysis, named ECG Joint Embedding Predictive Architecture (ECG-JEPA). ECG-JEPA employs a masking strategy to learn semantic representations of ECG data. Unlike existing methods, ECG-JEPA predicts at the hidden representation level rather than reconstructing raw data. This approach offers several advantages in the ECG domain: (1) it avoids producing unnecessary details, such as noise, which is common in standard ECG; and (2) it addresses the limitations of naïve L2 loss between raw signals. Another key contribution is the introduction of a special masked attention tailored for 12-lead ECG data, Cross-Pattern Attention (CroPA). CroPA enables the model to effectively capture inter-patch relationships. Additionally, ECG-JEPA is highly scalable, allowing efficient training on large datasets. Our code is openly available https://github.com/sehunfromdaegu/ECG_JEPA.
1 Introduction
Electrocardiograms (ECG) provide a non-invasive method to measure the electrical activity of the heart over time, serving as a crucial tool for diagnosing various cardiac conditions. While numerous supervised learning models have been developed to detect heart diseases using ECG data [1, 2, 3], these models often face significant performance degradation when applied to data distributions different from those on which they were trained.
Self-supervised learning (SSL) has emerged as a powerful paradigm for learning general representations across various domains, including natural language processing [4, 5, 6], computer vision [7, 8, 9], and video analysis [10, 11]. Despite its promise, applying SSL to ECG data presents unique challenges. For instance, data augmentation, which is essential in many SSL architectures, is more complex for ECG than for computer vision (CV) data. Simple transformations like rotation, scaling, and flipping, effective in CV, can distort the physiological meaning of ECG signals. Additionally, ECG recordings often contain artifacts and noise, which cause autoencoder-based SSL models to struggle with reconstructing raw signals. These architectures may also miss visually subtle but diagnostically critical features, such as P-waves and T-waves, which are imperative for diagnosing certain cardiac conditions.
In this work, we propose a novel ECG Joint-Embedding Predictive Architecture (ECG-JEPA) tailored for 12-lead ECG data, effectively addressing the aforementioned challenges. ECG-JEPA utilizes a transformer architecture to capture the semantic meanings of ECG. By masking several patches of the ECG data, ECG-JEPA predicts the abstract representations of the missing patches, indicating a high-level understanding of the data. Additionally, we develop a novel masked-attention for multi-lead ECG data, coined Cross-Pattern Attention (CroPA). CroPA enables the transformer architecture to effectively capture the relationships between patches in multi-lead ECG.
In this work, we demonstrate that:
  • ECG-JEPA significantly improves linear evaluation and fine-tuning on classification tasks compared to existing SSL methods without hand-crafted augmentations.
  • CroPA enhances the model’s ability to capture inter-lead relationships, leading to better performance on downstream tasks.
  • ECG-JEPA can also recover important ECG features, including heart rate and QRS duration, which are classical indicators used in ECG evaluation. This is the first work to demonstrate that learned representations can effectively recover various ECG features.
  • ECG-JEPA is highly scalable, allowing efficient training on large datasets. For instance, ECG-JEPA is trained for only 100 epochs, yet outperforms other ECG SSL models on most downstream tasks, taking approximately 26 hours on a single RTX 3090 GPU.
Figure 1: 12-lead ECG. The standard 12-lead ECG captures electrical activity from different angles around the heart, providing comprehensive cardiac assessment. Artifacts and noise, such as powerline interference and baseline wander, are common in ECG recordings. The figure shows an example of a 12-lead ECG with baseline wander.
2 Background
Self-Supervised Learning (SSL) facilitates learning abstract representations from input data without the need for labeled data, which is particularly beneficial in medical domains where labeled data is scarce and expensive. SSL leverages inherent data patterns to learn useful representations, allowing models to adapt to various downstream tasks with greater robustness to data imbalances [12].
2.1 Generative Architectures
Generative architectures involve reconstructing an input $x$ from its degraded version $x^{\prime}$ using an encoder-decoder framework. The idea behind generative architectures is that the ability to reconstruct the clean data from a corrupted one indicates a good understanding of the data. The encoder maps the perturbed input $x^{\prime}$ into a latent representation, which the decoder then uses to generate the original input $x$ [13].
2.2 Joint-Embedding Predictive Architectures
Joint-Embedding Predictive Architectures (JEPA) [14] process pairs $x$ and their corrupted versions $x^{\prime}$ to obtain representations $z$ and $z^{\prime}$ through encoders. Unlike generative architectures that predict in the input space, JEPA predicts in the hidden representation space by reconstructing $z$ from $z^{\prime}$. This approach effectively avoids the challenge of predicting unpredictable details, which is common in biological signals.
2.3 Electrocardiogram (ECG)
The electrocardiogram (ECG) is a non-invasive diagnostic method that records the heart’s electrical activity over time using electrodes placed on the skin. The standard 12-lead ECG captures electrical activity from multiple angles. These 12 leads are categorized into limb leads (I, II, III), augmented limb leads (aVR, aVL, aVF), and chest leads (V1-V6). Each lead provides unique information about the heart’s electrical activity, offering a comprehensive view that aids in diagnosing various cardiac conditions. Refer to Figure 1 for an illustration.
In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads. This choice maintains the necessary diagnostic information while optimizing computational efficiency.
Figure 2: Key ECG Features.
2.4 ECG Features
[…]
3 Methodology
ECG-JEPA is trained by predicting masked representations of ECG data within the hidden representation space, using only a partial view of the input. The proposed architecture utilizes a student-teacher network framework, as illustrated in Figure 3. We subdivide the multi-channel ECG into non-overlapping patches and sample a subset of these patches for masking. However, reconstructing the raw signals of masked patches can be particularly challenging in the ECG domain due to the prevalence of noise in biological signals. Instead, our model predicts the masked patches in the hidden representation space, where this challenge can be effectively addressed. We validate the quality of the learned representations through various downstream tasks, including linear probing, fine-tuning on classification tasks, and ECG feature extraction tasks.
3.1 Patch Masking
Let $x\in\mathbb{R}^{L\times T}$ represent a multi-lead ECG of length $T$ with $L$ channels. We divide the interval $[0,T]$ into $N$ non-overlapping subintervals of length $t$. Each subinterval in each channel constitutes a patch of $x$, resulting in $L\times N$ patches. The masking strategy in multi-lead ECG must be carefully chosen because patches in different leads at the same temporal position are highly correlated, potentially making the prediction task too easy. To address this, we mask all patches across different leads in the same temporal space. With this in mind, we employ two masking strategies: random masking and multi-block masking.
In random masking, we randomly select a percentage of subintervals to mask, while in multi-block masking, we select multiple consecutive subintervals to mask. Note that we allow these consecutive subintervals to overlap, which requires the model to predict much longer sequences of representations. In this paper, we use both masking strategies to evaluate the effectiveness of ECG-JEPA, with a random masking ratio of $(0.6,0.7)$ and a multi-block masking ratio of $(0.175,0.225)$ with a frequency of 4. The unmasked patches serve as the contextual input for the student networks, while the masked patches are the ones for which we aim to predict the representations.
The patches are then converted into sequences of token vectors using a linear layer, and augmented with positional embeddings. For positional encoding, we employ the conventional 2-dimensional sinusoidal positional embeddings for the student and teacher networks, while we use 1-dimensional sinusoidal positional embeddings for the predictor network.
Figure 3: ECG-JEPA. The student network processes only non-masked patches of time series, whereas the teacher network processes the entire time series. Then the predictor predicts the representation of masked patches processed by the teacher. Note that both student and teacher network process the multi-channel time series, while predictor gets only a single-channel time series. In this picture, there are $L=3$ leads with $N=5$ subintervals and $Q=3$ unmasked subintervals.
3.2 Teacher, Student, and Predictor
ECG-JEPA consists of three main components: the teacher network, the student network, and the predictor network. Both the teacher and student networks are based on standard transformer architectures. The weights of the teacher network are updated through an exponential moving average (EMA) of the student network, with details provided in the supplementary materials. The predictor network, a smaller transformer, operates on single-channel representations.
The teacher network handles the entire $L\times N$ patches, generating fully contextualized $L\times N$ representations. The student network, however, processes only $L\times Q$ visible (unmasked) patches, where $Q<N$ represents the number of visible time intervals. These $L\times Q$ representations from the student are then concatenated with the (learnable) mask tokens, resulting in $L\times N$ representations. Subsequently, each lead’s representations are passed to the predictor, which processes single-channel representations. The predictor’s output, the predicted representations of the target patches, is compared with the target representations using a smooth L1 loss function.
3.3 Cross-Pattern Attention (CroPA)
Biological signals often require simultaneous observation across multiple channels. In a 12-lead ECG, abnormalities must be consistent across multiple leads to diagnose specific cardiac pathologies. Therefore, the meaning of a single patch should be understood in the context of other patches within the same channel and across different channels. In contrast, patches in different channels and temporal spaces are often neglected when interpreting a 12-lead ECG.
To address this, we introduce Cross-Pattern Attention (CroPA), a masked self-attention tailored for multi-lead ECG data. CroPA allows each patch to attend only to patches within the same channel and temporal space (Figure 4). This method enhances our model’s effectiveness compared to the standard attention mechanism, where patches attend to all other patches indiscriminately.
4 Related Work
This section provides an overview of existing SSL methods that have been applied to ECG data. We compare ECG-JEPA with these methods on various downstream tasks in the next section.
4.1 Contrastive Learning
Contrastive learning has been widely explored in various research works for its effectiveness in unsupervised representation learning. Contrastive Predictive Coding (CPC) [15] aims to capture useful data representations by predicting future observations in the latent space. It uses sequential data where the “true future" is treated as the positive sample, and other observations within the same sequence are treated as negative samples. SimCLR [7] and MoCo [16] also leverage contrastive approaches, with positive samples being augmented versions of the same image and negative samples being other images. SimCLR relies on large batch sizes and extensive data augmentation, while MoCo uses a momentum encoder to maintain a dynamic dictionary of encoded samples, enabling more stable training with smaller batch sizes.
4.2 Generative Architectures
BERT [4] pioneered generative architectures in natural language processing (NLP), inspiring similar methods in other domains. Masked Autoencoders (MAE) [8] have been developed for computer vision, where a portion of the input image is masked, and the model is trained to reconstruct these missing parts. Recently, [17] pointed out that generative architectures often prioritize learning the principal subspaces of the data, which may limit their effectiveness in capturing semantic representations for perceptual tasks.
Figure 4: Cross-Pattern Attention (CroPA). The patch in the middle attends only to the colored patches, which are in the same lead and temporal space.
4.3 Joint-Embedding Predictive Architectures
Data2vec [18] introduced an SSL architecture that predicts the average of the encoder’s final layers’ outputs, which was further improved in efficiency in data2vec 2.0 [19]. Notable models like I-JEPA [9] and V-JEPA [11] predict the hidden representations of masked patches, enabling semantic representation learning for images and videos, respectively.
4.4 Self-Supervised Learning for ECG
Several attempts have been made to train SSL models for 12-lead ECG data. Contrastive Multi-segment Coding (CMSC) [20] is a contrastive method that divides an ECG into two segments, encouraging close representations for compatible segments while pushing apart incompatible ones. Contrastive Predictive Coding (CPC) [15] has also been used for ECG data in [21], predicting future representations of the ECG. A significant disadvantage of CPC is that it is not scalable to large datasets, as the LSTM modules inside CPC make the training process extremely slow. Recently, [22] utilized masked autoencoders for ECG data, proposing temporal and channel masking strategies, coined as Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE), respectively. Similarly, [23] proposed ST-MEM, an MAE for ECG data that masks random time intervals for each lead. One possible drawback of MLAE and ST-MEM is that predicting some patches can be relatively easy considering that 12 leads signals are highly correlated.
5 Experimental Settings
This section provides an overview of the experimental settings. Details of all experiments can be found in the supplementary materials.
5.1 Pretraining Datasets
Training SSL models with large datasets is crucial for developing generalized representations. However, most previous works have used relatively small datasets, with the exception of [23], where an SSL model was trained with a large number of 12-lead ECGs. Following [23], we use the Chapman [24], Ningbo [25], and CODE-15 [26] datasets for pretraining ECG-JEPA. The Chapman and Ningbo datasets collectively consist of 45,152 10-second 12-lead ECGs at 500Hz. CODE-15 includes 345,779 12-lead ECGs from 233,770 patients at 400Hz, with 143,328 being 10-second recordings. After excluding recordings with missing values, we have 43,240 ECGs from Chapman and Ningbo and 130,900 ECGs from CODE-15. All ECGs were resampled to 250Hz for consistency.
5.2 Downstream Datasets
We use the PTB-XL [27] and CPSC2018 [28] datasets to evaluate the performance of ECG-JEPA on downstream tasks. PTB-XL contains 21,837 clinical 10-second 12-lead ECG records from 18,885 patients, recorded at 500Hz and annotated with 71 diagnostic labels, which are aggregated into five superclasses. We use these superclass labels for our experiments. The CPSC2018 dataset includes 6,877 12-lead ECG recordings with nine annotated cardiac conditions. These datasets are multi-label in nature, where each recording can have multiple labels simultaneously. The details of the datasets are provided in the supplementary materials.
5.3 Pretraining
After resampling the ECG data to 250Hz, each 10-second multi-lead ECG has $T=2500$ time points. We divide the interval $[0,T]$ into $N=50$ non-overlapping subintervals of length $t=50$. The model is trained for 100 epochs without any data augmentations, and the final checkpoint is used for downstream tasks.
5.4 Architecture
We use vision transformer (ViT-Base) architectures [29] for the student and teacher network, and a smaller standard transformer architecture for the predictor network. The student and teacher network process the multi-lead ECG, while the predictor processes each lead independently to reconstruct the masked representations. See supplementary materials for detailed architecture specifications.
5.5 Downstream Tasks
We conduct extensive experiments to show that ECG-JEPA effectively captures semantic representations. We evaluate its performance on classification tasks through linear probing and fine-tuning. Additionally, we evaluate the performance on the low-shot setting, and on the reduced-lead setting where the downstream dataset is restricted to have a single or two leads.
Next, we validate the expressive power of the learned representations by predicting ECG features, including heart rate and QRS duration. To the best of our knowledge, this is the first work showing that the learned representations can recover various ECG features. The successful prediction of these features provides compelling evidence that the representations are not only informative but also capture essential and clinically relevant characteristics, making them highly reliable for ECG analysis.
In ECG analysis, it is common for a single recording to have multiple labels simultaneously, making datasets like PTB-XL and CPSC2018 multi-label tasks. However, many previous studies have converted this into a multi-class classification problem by restricting the dataset to subsets with only a single label. To ensure a fair comparison, we will pretrain the methods using publicly available code and evaluate their performance on the multi-label classification task. If the code is not available, we will convert our task into a multi-class problem and compare the results with the reported performance in the literature.
6 Experiments
In this section, we assess the performance of the learned representations on various downstream tasks. Our experiments aim to demonstrate that the learned representations are generalizable and effective in capturing essential ECG features. We compare ECG-JEPA with several state-of-the-art SSL methods.
For classification tasks, we use AUC (Area Under the ROC Curve) and F1 scores as our primary evaluation metrics. AUC evaluates model performance across all possible thresholds, providing a comprehensive and stable measure of discriminative ability. In contrast, the F1 score balances precision and recall at a fixed threshold. In our experiments, we use the AUC as our main metric, as it provides a more robust evaluation of the model’s performance across varying classification thresholds, ensuring that the results are not overly dependent on a single decision boundary.
For multi-label classification, the AUC is computed by averaging the scores of binary classifications for each label. For multi-class classification, the AUC is calculated using the average of one-vs-rest approach. In both cases, the F1 score is computed using a macro-averaged approach that averages the F1 scores of all classes.
In most cases, ECG-JEPA outperforms other SSL methods that leverage hand-crafted augmentations, demonstrating its effectiveness in learning general representations of ECG data. In experiments, E-JEPArb and E-JEPAmb refer to ECG-JEPA models trained with random masking and multi-block masking strategies, respectively.
6.1 Linear Evaluation
Tables 1 and 2 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets. We train a linear classifier on top of the frozen representations for 10 epochs, and evaluate its performance on downstream tasks. As shown in the tables, ECG-JEPA consistently outperforms other SSL methods, demonstrating superior efficiency and effectiveness with substantially reduced computational resources.
Table 1: Linear evaluation on multi-label task.
PTB-XL CPSC2018
Method Epochs AUC F1 AUC F1
ST-MEM 800 0.896 0.662 0.964 0.752
SimCLR 300 0.866 0.624 0.890 0.523
CMSC 300 0.802 0.472 0.767 0.206
CPC 100 0.620 0.167 0.687 0.091
E-JEPArb 100 0.906 0.690 0.969 0.769
E-JEPAmb 100 0.912 0.712 0.971 0.789
Table 2: Linear evaluation on multi-class task.
PTB-XL CPSC2018
Method Epochs AUC F1 AUC F1
MoCo v3111Score as reported in [23]. 800 0.739 0.142 0.712 0.080
MTAE [^1] Score as reported in [23]. 800 0.807 0.437 0.818 0.349
MLAE [^1] Score as reported in [23]. 800 0.779 0.382 0.794 0.263
ST-MEM 800 0.888 0.566 0.973 0.805
SimCLR 300 0.842 0.496 0.918 0.624
CMSC 300 0.796 0.442 0.787 0.391
CPC 100 0.600 0.201 0.672 0.210
E-JEPArb 100 0.894 0.616 0.974 0.805
E-JEPAmb 100 0.896 0.628 0.973 0.819
6.2 Fine-tuning
Fine-tuning is another method to evaluate the quality of learned representations, as it tests the model’s ability to adapt its pre-trained features to new tasks. In the fine-tuning process, we add a linear classification head at the end of the encoder and train the whole network for 10 epochs. Fine-tuning can potentially enhance performance beyond what is achieved with linear evaluation alone.
Table 3 presents the results of fine-tuning on the PTB-XL and CPSC2018 datasets. ECG-JEPA is compared with other SSL methods in a multi-class classification setting. The results indicate that ECG-JEPA achieves the highest AUC and F1 scores on PTB-XL and the highest AUC on CPSC2018. These results highlight ECG-JEPA’s effectiveness in fine-tuning scenarios, demonstrating its ability to leverage transfer learning by refining and adapting its learned representations for improved performance in specific classification tasks.
Table 3: Fine-tuning on multi-class task.
PTB-XL CPSC2018
Method Epochs AUC F1 AUC F1
MoCo v3111Score as reported in [23]. 800 0.913 0.644 0.967 0.838
MTAE [^1] Score as reported in [23]. 800 0.910 0.613 0.961 0.769
MLAE [^1] Score as reported in [23]. 800 0.915 0.625 0.973 0.816
CMSC [^1] Score as reported in [23]. 800 0.877 0.510 0.938 0.717
ST-MEM 800 0.929 0.668 0.977 0.820
SimCLR 300 0.905 0.650 0.934 0.693
CPC [^2] We did not fine-tune CPC due to its slow training process. 100 - - - -
E-JEPArb 100 0.944 0.710 0.980 0.821
E-JEPAmb 100 0.937 0.680 0.983 0.799
6.3 Low-shot Linear Evaluation
Table 4 presents the performance comparison on the low-shot task. Low-shot learning is particularly challenging, as models must generalize effectively with limited labeled data. Given the difficulty and resource-intensive nature of obtaining labeled data in medical research, low-shot learning represents a realistic and critical scenario in the medical field. In this experiment, we evaluate the performance of ECG-SSL models on the PTB-XL multi-label task with only 1% and 10% of the training set, while keeping the test set fixed. As shown in the table, ECG-JEPA demonstrates a clear advantage over other SSL methods, with its effectiveness becoming particularly evident in low-shot learning tasks. This suggests that ECG-JEPA can be particularly well-suited for transfer learning where labeled data is scarce.
Table 4: Low-shot linear evaluation on the multi-label PTB-XL task. The mean and standard deviation of macro AUC are reported for 1% (192 samples) and 10% (1923 samples) of the training set, selected three times independently.
PTB-XL
Method Epochs 1% 10%
ST-MEM 800 0.807 $\pm$ 0.005 0.872 $\pm$ 0.001
SimCLR 300 0.803 $\pm$ 0.002 0.843 $\pm$ 0.001
CMSC 300 0.750 $\pm$ 0.008 0.792 $\pm$ 0.001
CPC 100 0.523 ±$\pm 0.006$ 0.560 $\pm$ 0.005
E-JEPArb 100 0.836 $\pm$ 0.006 0.887 $\pm$ 0.000
E-JEPAmb 100 0.843 $\pm$ 0.004 0.894 $\pm$ 0.003
6.4 Reduced Lead Evaluation
Because transformer architectures can accept variable input lengths, we can evaluate ECG-JEPA’s performance on reduced lead tasks. In this experiment, we perform a linear evaluation on the PTB-XL multi-label task using reduced leads. Specifically, we utilize only a single lead (Lead II) and two leads (Lead II and V1), training linear classifiers on top of the learned representations for 10 epochs. We then compare ECG-JEPA’s performance with ST-MEM, another model supporting reduced lead evaluation. Table 5 presents the results of this reduced lead linear evaluation. Notably, ECG-JEPA performs well even in the reduced lead setting. This robustness is particularly valuable since most daily mobile devices typically output only one or two leads, highlighting ECG-JEPA’s potential for practical applications in mobile health monitoring.
Table 5: Reduced lead evaluation. Linear evaluation on PTB-XL multi-label task with reduced lead settings.
1-Lead 2-Lead
Method AUC F1 AUC F1
ST-MEM 0.832 0.571 0.848 0.597
E-JEPArb 0.846 0.596 0.877 0.647
E-JEPAmb 0.849 0.593 0.880 0.657
6.5 ECG Feature Extraction
ECG features are essential for diagnosing and monitoring cardiac conditions. We evaluate the model’s ability to extract key ECG features, including heart rate and QRS duration, from the learned representations of the PTB-XL dataset. Unlike classification tasks, which are highly perceptual, ECG features are directly related to the morphology of the signal. For example, heart rate can be measured by the average distance between R-peaks, and QRS duration can be measured by the average width of the QRS complex.
Several methods have been proposed to segment ECG signals [30, 31, 32, 33]. Segmentation models can be used to extract ECG features. In this experiment, we employed a publicly available segmentation model [33] to extract ECG features from the PTB-XL dataset, which served as our ground truth labels. We then trained a linear regression model on the learned representations to predict these ECG features, using the mean squared error (MSE) as the loss function.
Table 6 presents the performance comparison, showing the means and standard deviations of the absolute differences between the extracted and predicted values across all samples in the PTB-XL test set.
Interestingly, although the model’s learned representations are implied to capture high-level features based on their strong performance in linear evaluation tasks across various datasets, they still retain the ability to recover low-level information, such as ECG features. This dual capability of capturing both high-level and low-level information is noteworthy and highlights the model’s versatility.
Table 6: ECG feature extraction. Means and standard deviations of absolute differences between the predicted and extracted values for heart rate and QRS duration. Small values indicate better performance. The final row provides the overall mean and standard deviation for both features across all samples in the PTB-XL test set.
Method Heart Rate (BPM) QRS Dur. (ms)
ST-MEM 1.35 $\pm$ 2.38 4.60 $\pm$ 4.16
SimCLR 1.87 $\pm$ 2.81 6.14 $\pm$ 5.80
CMSC 7.20 $\pm$ 7.43 10.12 $\pm$ 9.98
CPC 11.40 $\pm$ 11.04 11.55 $\pm$ 11.55
E-JEPArb 1.54 $\pm$ 2.62 4.81 $\pm$ 4.29
E-JEPAmb 1.45 $\pm$ 2.44 4.41 $\pm$ 4.08
Mean 75.01 $\pm$ 17.65 90.48 $\pm$ 17.02
7 Ablation Study
7.1 Effect of CroPA
Table 7 presents the results of our evaluation of the effectiveness of CroPA. CroPA introduces a “human-like" inductive bias, enabling the model to be trained more efficiently on multi-lead ECG data. Without CroPA, models may require more epochs to converge. For a fair comparison, we trained ECG-JEPA with and without CroPA for 100 and 200 epochs and compared their performance on the PTB-XL multi-class task. The results show that CroPA considerably improves the model’s performance, demonstrating its effectiveness in capturing inter-lead relationships and enhancing the model’s ability to learn meaningful representations.
Table 7: Effect of CroPA. Linear evaluation (lin) and fine-tuning (ft) results on PTB-XL multi-class task with ECG-JEPArb and ECG-JEPAmb. In both cases, CroPA improves the model’s performance.
lin ft
Mask CroPA Epochs AUC AUC
Random x 100 0.888 0.930
Random x 200 0.887 0.927
Random o 100 0.894 0.944
Multi-block x 100 0.872 0.924
Multi-block x 200 0.886 0.914
Multi-block o 100 0.896 0.937
7.2 Masking Ratio
Table 8 presents the performance of ECG-JEPA in linear evaluation with different masking ratios and strategies. The results indicate that the model benefits from a high masking ratio. Notably, multi-block masking is advantageous for linear evaluation, while random masking is more effective for fine-tuning, as indicated in Table 3. Although random masking with a ratio of (0.7, 0.8) achieves better performance in the PTB-XL multi-label task, a masking ratio of (0.6, 0.7) performs better in other tasks. Therefore, we chose the latter for our main experiments.
Table 8: Effect of masking strategy. Linear evaluation results on PTB-XL multi-label task using different masking ratios and strategies.
Mask Ratio Freq. AUC F1
Random (0.3, 0.4) 1 0.884 0.652
Random (0.4, 0.5) 1 0.904 0.698
Random (0.5, 0.6) 1 0.906 0.697
Random (0.6, 0.7) 1 0.906 0.690
Random (0.7, 0.8) 1 0.909 0.706
Multi-block (0.10, 0.15) 4 0.904 0.678
Multi-block (0.15, 0.20) 4 0.905 0.687
Multi-block (0.175, 0.225) 4 0.912 0.712
8 Future Work
Our method has potential applications in various physiological multivariate signals, such as EEG and EMG. These signals share characteristics with ECG, including their multivariate nature and high dimensionality, making them suitable for our proposed approach.
Another promising direction involves the integration of multi-modal data. For example, multi-lead ECG data could be combined with other diagnostic modalities, such as chest X-rays (CXR), to provide a more comprehensive understanding of a patient’s condition.
One significant challenge in pursuing these extensions is the scarcity of large-scale datasets. Addressing this limitation is crucial for advancing and validating our method across diverse applications.
9 Conclusion
We proposed ECG-JEPA, a novel SSL method tailored for 12-lead ECG data. By utilizing a JEPA coupled with the innovative relative positional encoding method, CroPA, ECG-JEPA effectively learns meaningful representations of ECG signals. This approach addresses the challenges posed by noise and artifacts in ECG data, demonstrating substantial improvements over existing SSL methods in various downstream tasks, with the added benefit of significantly faster convergence.
Our extensive experimental evaluations reveal that ECG-JEPA outperforms state-of-the-art SSL methods across several tasks, including linear evaluation, fine-tuning, low-shot learning, and ECG feature extraction. Moreover, our investigation into the use of 8 leads, as opposed to the full 12-lead ECG, indicates that this reduction does not compromise performance while optimizing computational efficiency. This finding is particularly significant for applications constrained by limited computational resources.
Appendix A Supplementary Materials
A.1 Experimental Details on ECG-JEPA
Hyperparameters for ECG-JEPA pretraining, linear evaluation, and fine-tuning are provided in Tables 9, 10, and 11, respectively. In ECG-JEPAmb, the number of visible patches in ECG-JEPAmb varies more than in ECG-JEPArb, resulting in higher GPU memory usage. Consequently, we reduced the batch size to 64 to fit the model on a single NVIDIA RTX 3090 GPU. Interestingly, ECG-JEPAmb benefits from larger learning rates, even with the halved batch size.
For fine-tuning process, the actual learning rate is calculated as $lr=base\text{\_}lr\times batchsize/256$, following the heuristic by [34].
Table 9: Pretraining Settings for ECG-JEPA.
[…]
optimizer AdamW AdamW
learning rate 2.5e-5 5e-5
weight decay 0.05 0.05
batch size 128 64
learning rate schedule cosine decay cosine decay
warmup epochs 5 5
epochs 100 100
drop path 0.1 0.1
Table 10: Linear Evaluation Settings
config value
optimizer AdamW
learning rate 5e-4
weight decay 0.05
batch size 32
learning rate schedule cosine decay
warmup epochs 3
epochs 10
Table 11: Fine-tuning Settings.
config value
optimizer AdamW
base learning rate 1.0e-4
weight decay 0.05
batch size 16
learning rate schedule cosine decay
warmup epochs 3
epochs 10
A.2 Pretraining Other SSL Models
Besides pretraining ECG-JEPA, we also pretrained other models, including CMSC [20], CPC [15], and SimCLR [7] using the same datasets as ECG-JEPA.
For CMSC and CPC, we adhered to the original architecture and hyperparameters. SimCLR utilized a ResNet50 [35] encoder with an output dimension of 2048. CMSC and SimCLR were pretrained for 300 epochs, selecting the best checkpoint at 100, 200, or 300 epochs based on linear evaluation performance on the PTB-XL multi-label setting. Due to the slow training process, CPC was pretrained for only 100 epochs, taking approximately 9 days on a single NVIDIA RTX 3090 GPU due to the LSTM module in the model. For ST-MEM [23], we employed the publicly available checkpoint pretrained for 800 epochs.
Given SimCLR’s sensitivity to data augmentations, we applied several that work well empirically: baseline shift (adding a constant to all leads), baseline wander (low-frequency noise), Gaussian noise (random noise), powerline noise (50Hz noise), channel resize, random crop, and jump noise (sudden jumps). These augmentations aimed to enhance the robustness of the model to various signal distortions.
A.3 Downstream Datasets Details
Table 12, 13, 14, and 15 show the distribution of the PTB-XL and CPSC2018 datasets in both multi-label and multi-class settings.
The PTB-XL dataset is stratified into ten folds, where the first eight folds are used for training, the ninth fold for validation, and the tenth fold for testing. In our experiments, we used the first nine folds for training and the tenth fold for testing, as we did not observe overfitting during linear evaluation and fine-tuning.
For the CPSC2018 dataset, only the training set is publicly available, which is stratified into seven folds. We used the first six folds for training and the seventh fold for testing, omitting the validation set. The original CPSC2018 dataset consists of 6,877 ECG recordings, but we excluded recordings with a length of less than 10 seconds, resulting in 6,867 ECG recordings.
Table 12: PTB-XL multi-label distribution. The sum of samples in each class exceeds the total number of ECG recordings due to the multi-label nature of the dataset.
# ECG Norm MI STTC CD HYP
Multi-label 21799 9514 5469 5235 4898 2649
Train 19230 8551 4919 4714 4402 2387
Test 2158 963 550 521 496 262
Table 13: PTB-XL multi-class distribution. ECGs with multiple labels are excluded from the PTB-XL multi-label dataset.
# ECG Norm MI STTC CD HYP
Multi-class 16244 9069 2532 2400 1708 535
Train 14594 8157 2276 2158 1524 479
Test 1650 912 256 242 184 56
Table 14: CPSC2018 multi-label distribution. The sum of samples in each class exceeds the total number of ECG recordings due to the multi-label nature of the dataset.
# ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
Multi-label 6867 918 1220 235 220 721 614 699 868 1854
Train 5989 805 1059 206 197 632 534 615 742 1616
Test 878 113 161 29 23 89 80 84 126 238
Table 15: CPSC2018 multi-class distribution. The number of ECG recordings for each superclass in the PTB-XL dataset. Note that the sum of samples in each class exceeds the total number of ECG recordings due to multi-label the nature of the dataset.
# ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
Multi-class 6391 918 975 178 185 685 531 606 783 1530
Train 5577 805 849 159 169 600 459 534 671 1331
Test 814 113 126 19 16 85 72 72 112 199
A.4 Exponential Moving Average
The teacher network is initialized as a copy of the student network and is updated using an exponential moving average (EMA) of the student’s weights. The EMA is computed as follows:
$\theta_{\text{teacher}}^{i}=\beta_{i}\theta_{\text{teacher}}^{i-1}+(1-\beta_{i})\theta_{\text{student}}^{i}$
where $i$ denotes the current training iteration, and $\beta_{i}$ is a momentum parameter that evolves during training. The momentum parameter $\beta_{i}$ is computed as:
$\beta_{i}=\text{ema}_{0}+\frac{i\cdot(\text{ema}_{1}-\text{ema}_{0})}{\text{iterations\_per\_epoch}\cdot\text{epochs}}$
Here, $\text{ema}_{0}$ and $\text{ema}_{1}$ represent the initial and final values of the momentum parameter, respectively. For our implementation, $\text{ema}_{0}=0.996$ and $\text{ema}_{1}=1.0$.
A.5 Sufficiency of 8 Leads in ECG Analysis
In a standard 12-lead ECG, leads III, aVR, aVL, and aVF can be derived from leads I and II, as shown by the following equations:
[…]
$\text{aVR}=-\frac{(\text{I}+\text{II})}{2}$
$\text{aVL}=\text{I}-\frac{\text{II}}{2}$
$\text{aVF}=\text{II}-\frac{\text{I}}{2}$
These derivations demonstrate that leads III, aVR, aVL, and aVF do not provide additional independent information beyond what is captured by leads I and II. Therefore, using only 8 leads (I, II, V1, V2, V3, V4, V5, and V6) can be sufficient for effective ECG analysis. This simplification not only reduces the computational load but also maintains the diagnostic integrity of the analysis.
A.6 Comparison with 12-Lead Model
We now investigate the practical sufficiency of using 8 leads for ECG-JEPA pretraining. To evaluate the impact of this reduction, we trained models using both 8 leads and 12 leads and compared their performance on the linear evaluation of a multi-label task for PTB-XL.
Table 16 presents the results of this comparison using ECG-JEPArb. As expected, the performance difference between the 8-lead and 12-lead models is minimal, indicating that using 8 leads is sufficient for effective pretraining without significant loss of information.
Table 16: Comparison of 8-Lead and 12-Lead Models.
Model epochs AUC F1
8-Lead 100 0.906 0.690
12-Lead 100 0.905 0.699

### v2

@@
 Additionally, ECG-JEPA is highly scalable, allowing efficient training on large datasets.
-Our code is openly available https://github.com/sehunfromdaegu/ECG_JEPA.
+Our code is openly available at https://github.com/sehunfromdaegu/ECG_JEPA.
 1 Introduction
 Electrocardiograms (ECG) provide a non-invasive method to measure the electrical activity of the heart over time, serving as a crucial tool for diagnosing various cardiac conditions.
-While numerous supervised learning models have been developed to detect heart diseases using ECG data [1, 2, 3], these models often face significant performance degradation when applied to data distributions different from those on which they were trained.
-Self-supervised learning (SSL) has emerged as a powerful paradigm for learning general representations across various domains, including natural language processing [4, 5, 6], computer vision [7, 8, 9], and video analysis [10, 11].
+While numerous supervised methods have been developed to detect heart diseases using ECG data [1, 2, 3], these models often face significant performance degradation when applied to data distributions different from those on which they were trained.
+Self-supervised learning (SSL) has emerged as a powerful paradigm for learning general representations across various domains, including natural language processing (NLP) [4, 5, 6], computer vision (CV) [7, 8, 9], and video analysis [10, 11].
 Despite its promise, applying SSL to ECG data presents unique challenges.
-For instance, data augmentation, which is essential in many SSL architectures, is more complex for ECG than for computer vision (CV) data.
+For instance, data augmentation, which is essential in many SSL architectures, is more complex for ECG than for computer vision data.
 Simple transformations like rotation, scaling, and flipping, effective in CV, can distort the physiological meaning of ECG signals.
@@
 Additionally, we develop a novel masked-attention for multi-lead ECG data, coined Cross-Pattern Attention (CroPA).
-CroPA enables the transformer architecture to effectively capture the relationships between patches in multi-lead ECG.
-In this work, we demonstrate that:
-  • ECG-JEPA significantly improves linear evaluation and fine-tuning on classification tasks compared to existing SSL methods without hand-crafted augmentations.
-  • CroPA enhances the model’s ability to capture inter-lead relationships, leading to better performance on downstream tasks.
+CroPA incorporates clinical knowledge into the model as an inductive bias, guiding it to focus on clinically relevant patterns and relationships across leads.
+We demonstrate the following contributions of ECG-JEPA:
+  • ECG-JEPA achieves notable improvements in linear evaluation and fine-tuning on classification tasks compared to existing SSL methods without hand-crafted augmentations.
+  • CroPA introduces a specialized masked attention mechanism, allowing the model to focus on clinically relevant information in multi-lead ECG data, resulting in improved downstream task performance.
   • ECG-JEPA can also recover important ECG features, including heart rate and QRS duration, which are classical indicators used in ECG evaluation.
@@
 SSL leverages inherent data patterns to learn useful representations, allowing models to adapt to various downstream tasks with greater robustness to data imbalances [12].
+In this section, we briefly explain key SSL techniques in Sections 2.1 and 2.2, including generative architectures and joint-embedding predictive architectures.
+Additionally, we provide an overview of the ECG and its key features in Sections 2.3 and 2.4, highlighting the critical characteristics essential for understanding ECG data.
 2.1 Generative Architectures
@@
 Refer to Figure 1 for an illustration.
-In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads.
+In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads (see C).
 This choice maintains the necessary diagnostic information while optimizing computational efficiency.
@@
 […]
+Figure 3: ECG-JEPA.
+The teacher network processes the complete ECG data, generating fully contextualized representations.
+In contrast, the student network receives only the non-masked patches.
+The student’s representations are then passed to the predictor, which reconstructs the fully contextualized representations produced by the teacher.
+The predicted 6 representations are then compared with the corresponding teacher’s representation.
+This figure illustrates $L=3$ leads with $N=5$ subintervals and $Q=3$ unmasked subintervals.
 3 Methodology
-ECG-JEPA is trained by predicting masked representations of ECG data within the hidden representation space, using only a partial view of the input.
-The proposed architecture utilizes a student-teacher network framework, as illustrated in Figure 3.
+ECG-JEPA is trained by predicting masked representations of ECG data in the hidden representation space, using only a partial view of the input.
+The proposed architecture utilizes a student-teacher framework, as illustrated in Figure 3.
 We subdivide the multi-channel ECG into non-overlapping patches and sample a subset of these patches for masking.
@@
 The unmasked patches serve as the contextual input for the student networks, while the masked patches are the ones for which we aim to predict the representations.
-The patches are then converted into sequences of token vectors using a linear layer, and augmented with positional embeddings.
-For positional encoding, we employ the conventional 2-dimensional sinusoidal positional embeddings for the student and teacher networks, while we use 1-dimensional sinusoidal positional embeddings for the predictor network.
-Figure 3: ECG-JEPA.
-The student network processes only non-masked patches of time series, whereas the teacher network processes the entire time series.
-Then the predictor predicts the representation of masked patches processed by the teacher.
-Note that both student and teacher network process the multi-channel time series, while predictor gets only a single-channel time series.
-In this picture, there are $L=3$ leads with $N=5$ subintervals and $Q=3$ unmasked subintervals.
+The patches are converted into sequences of token vectors using a linear layer, and augmented with positional embeddings.
+We employ the conventional 2-dimensional sinusoidal positional embeddings for the student and teacher networks, while we use 1-dimensional sinusoidal positional embeddings for the predictor network.
 3.2 Teacher, Student, and Predictor
@@
 Both the teacher and student networks are based on standard transformer architectures.
-The weights of the teacher network are updated through an exponential moving average (EMA) of the student network, with details provided in the supplementary materials.
-The predictor network, a smaller transformer, operates on single-channel representations.
+The weights of the teacher network are updated using an exponential moving average (EMA) of the student network, as detailed in B.
+The predictor network, a smaller transformer, operates on single-channel representations, which still encode information from all leads due to the self-attention mechanism.
 The teacher network handles the entire $L\times N$ patches, generating fully contextualized $L\times N$ representations.
@@
 3.3 Cross-Pattern Attention (CroPA)
-Biological signals often require simultaneous observation across multiple channels.
-In a 12-lead ECG, abnormalities must be consistent across multiple leads to diagnose specific cardiac pathologies.
-Therefore, the meaning of a single patch should be understood in the context of other patches within the same channel and across different channels.
-In contrast, patches in different channels and temporal spaces are often neglected when interpreting a 12-lead ECG.
-To address this, we introduce Cross-Pattern Attention (CroPA), a masked self-attention tailored for multi-lead ECG data.
-CroPA allows each patch to attend only to patches within the same channel and temporal space (Figure 4).
-This method enhances our model’s effectiveness compared to the standard attention mechanism, where patches attend to all other patches indiscriminately.
+Multi-lead ECG signals require careful analysis of patterns that are often consistent across different leads, which is crucial for identifying potential cardiac abnormalities.
+This demands attention mechanisms that prioritize relationships within the same lead and within relevant time windows.
+To incorporate this structural insight, we introduce Cross-Pattern Attention (CroPA), a masked self-attention mechanism designed for multi-lead ECG data.
+CroPA imposes an inductive bias by allowing each patch to attend only to patches within the same lead and temporal space (Figure 4).
+This aligns with the way ECG signals are typically interpreted, where intra-lead and temporally adjacent signals hold the most significance.
+By incorporating this inductive bias, CroPA helps the model focus on relevant intra-lead relationships, reducing interference from unrelated signals across different channels and time points.
+Compared to the standard self-attention mechanism, which treat all patches equally, CroPA reflects a structured approach that mirrors the process of multi-lead signal interpretation, leading to improved performance in downstream tasks.
 4 Related Work
-This section provides an overview of existing SSL methods that have been applied to ECG data. We compare ECG-JEPA with these methods on various downstream tasks in the next section.
+This section provides an overview of existing SSL methods that have been applied to ECG data. We compare ECG-JEPA with these methods on various downstream tasks in Section 6.
 4.1 Contrastive Learning
@@
 4.2 Generative Architectures
-BERT [4] pioneered generative architectures in natural language processing (NLP), inspiring similar methods in other domains.
-Masked Autoencoders (MAE) [8] have been developed for computer vision, where a portion of the input image is masked, and the model is trained to reconstruct these missing parts.
-Recently, [17] pointed out that generative architectures often prioritize learning the principal subspaces of the data, which may limit their effectiveness in capturing semantic representations for perceptual tasks.
+Generative architectures, first popularized by BERT [4] in natural language processing (NLP), have since inspired a variety of approaches in other domains.
+For example, in computer vision, Masked Autoencoders (MAE) [8] extend this concept by masking portions of an input image and training the model to reconstruct the missing parts, similar to how BERT predicts masked tokens in text.
+Recently, the authors of [17] pointed out that generative architectures often prioritize learning the principal subspaces of the data, which may limit their effectiveness in capturing semantic representations for perceptual tasks.
 Figure 4: Cross-Pattern Attention (CroPA). The patch in the middle attends only to the colored patches, which are in the same lead and temporal space.
 4.3 Joint-Embedding Predictive Architectures
-Data2vec [18] introduced an SSL architecture that predicts the average of the encoder’s final layers’ outputs, which was further improved in efficiency in data2vec 2.0 [19].
-Notable models like I-JEPA [9] and V-JEPA [11] predict the hidden representations of masked patches, enabling semantic representation learning for images and videos, respectively.
+In contrast to generative approaches, which focus on reconstructing data in the input space, joint-embedding predictive architectures (JEPA) aim to predict in the representation space, allowing models to learn more abstract and semantic features.
+An example of this approach is Data2vec [18], which introduced an SSL architecture that predicts the average of the encoder’s final layer outputs.
+Other examples include I-JEPA [9] and V-JEPA [11], which predict the hidden representations of masked patches, enabling semantic representation learning for images and videos, respectively.
 4.4 Self-Supervised Learning for ECG
-Several attempts have been made to train SSL models for 12-lead ECG data.
-Contrastive Multi-segment Coding (CMSC) [20] is a contrastive method that divides an ECG into two segments, encouraging close representations for compatible segments while pushing apart incompatible ones.
-Contrastive Predictive Coding (CPC) [15] has also been used for ECG data in [21], predicting future representations of the ECG.
-A significant disadvantage of CPC is that it is not scalable to large datasets, as the LSTM modules inside CPC make the training process extremely slow.
-Recently, [22] utilized masked autoencoders for ECG data, proposing temporal and channel masking strategies, coined as Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE), respectively.
-Similarly, [23] proposed ST-MEM, an MAE for ECG data that masks random time intervals for each lead.
-One possible drawback of MLAE and ST-MEM is that predicting some patches can be relatively easy considering that 12 leads signals are highly correlated.
+Various SSL models have been explored for 12-lead ECG data.
+Contrastive Multi-segment Coding (CMSC) [19] splits an ECG into two segments, encouraging similar representations for compatible segments while separating incompatible ones.
+Contrastive Predictive Coding (CPC) [15], applied in [20], predicts future ECG representations, but its reliance on LSTM modules makes it inefficient for large datasets.
+More recently, [21] introduced masked autoencoders for ECG, proposing temporal and channel masking strategies, Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE).
+Similarly, [22] proposed ST-MEM, which masks random time intervals for each lead.
+However, both MLAE and ST-MEM may struggle with the high correlations between ECG leads, making some predictions too easy.
 5 Experimental Settings
-This section provides an overview of the experimental settings. Details of all experiments can be found in the supplementary materials.
+In all experiments, 10-second multi-lead ECG signals were resampled to 250Hz, yielding $T=2500$ time points.
+We divided the interval $[0,T]$ into $N=50$ non-overlapping subintervals, each of length $t=50$.
+The model was trained for 100 epochs without data augmentation, and the final checkpoint was used for downstream tasks.
+Additional experimental details are provided in Appendix A.
 5.1 Pretraining Datasets
@@
 After excluding recordings with missing values, we have 43,240 ECGs from Chapman and Ningbo and 130,900 ECGs from CODE-15.
-All ECGs were resampled to 250Hz for consistency.
 5.2 Downstream Datasets
@@
 These datasets are multi-label in nature, where each recording can have multiple labels simultaneously.
-The details of the datasets are provided in the supplementary materials.
-5.3 Pretraining
-After resampling the ECG data to 250Hz, each 10-second multi-lead ECG has $T=2500$ time points.
-We divide the interval $[0,T]$ into $N=50$ non-overlapping subintervals of length $t=50$.
-The model is trained for 100 epochs without any data augmentations, and the final checkpoint is used for downstream tasks.
+The details of the datasets are provided in Appendix A.3.
 5.3 Architecture
-We use vision transformer (ViT-Base) architectures [29] for the student and teacher network, and a smaller standard transformer architecture for the predictor network.
-The student and teacher network process the multi-lead ECG, while the predictor processes each lead independently to reconstruct the masked representations.
-See supplementary materials for detailed architecture specifications.
+Our model employs transformer encoder architectures for the student, teacher, and predictor networks.
+Both the teacher and student networks consist of 12 layers with 16 attention heads and a hidden dimension of 768.
+The predictor network, designed as a smaller transformer encoder, comprises 6 layers with 12 attention heads and a hidden dimension of 384.
+While the teacher and student networks process the multi-lead ECG data holistically, the predictor operates on each lead independently to reconstruct the masked representations.
+Importantly, this does not imply that the predictor relies solely on single-lead information for the reconstruction task; due to the self-attention mechanism, the input representations for each lead still encapsulate information from all leads.
 5.4 Downstream Tasks
 We conduct extensive experiments to show that ECG-JEPA effectively captures semantic representations.
-We evaluate its performance on classification tasks through linear probing and fine-tuning.
-Additionally, we evaluate the performance on the low-shot setting, and on the reduced-lead setting where the downstream dataset is restricted to have a single or two leads.
-Next, we validate the expressive power of the learned representations by predicting ECG features, including heart rate and QRS duration.
-To the best of our knowledge, this is the first work showing that the learned representations can recover various ECG features.
-The successful prediction of these features provides compelling evidence that the representations are not only informative but also capture essential and clinically relevant characteristics, making them highly reliable for ECG analysis.
-In ECG analysis, it is common for a single recording to have multiple labels simultaneously, making datasets like PTB-XL and CPSC2018 multi-label tasks.
-However, many previous studies have converted this into a multi-class classification problem by restricting the dataset to subsets with only a single label.
-To ensure a fair comparison, we will pretrain the methods using publicly available code and evaluate their performance on the multi-label classification task.
-If the code is not available, we will convert our task into a multi-class problem and compare the results with the reported performance in the literature.
+Its performance is evaluated on classification tasks using linear probing and fine-tuning.
+Furthermore, we assess its capability in low-shot learning settings, as well as under reduced-lead conditions where the downstream dataset is limited to single or two leads.
+Reduced-lead configurations are common in clinical practice, especially in scenarios like wearable devices or remote monitoring, where using the full 12-lead ECG setup is impractical.
+To validate the expressiveness of the learned representations, we predict key ECG features such as heart rate and QRS duration.
+Notably, this work is the first to show that these learned representations can recover a variety of ECG features.
+The ability to predict these features highlights the informativeness of the representations and their potential to capture clinically relevant characteristics, which is crucial for reliable ECG analysis.
+ECG datasets, such as PTB-XL and CPSC2018, often include multiple simultaneous labels for a single recording, making them multi-label tasks.
+However, many prior studies have simplified this into a multi-class classification problem by focusing on single-label subsets of the data.
+To ensure a fair comparison, we pretrain competing methods using publicly available code and evaluate them on the multi-label classification task.
+In cases where the code is unavailable, we will convert our task into a multi-class problem to align with the reported performance in the literature.
 6 Experiments
-In this section, we assess the performance of the learned representations on various downstream tasks.
-Our experiments aim to demonstrate that the learned representations are generalizable and effective in capturing essential ECG features.
-We compare ECG-JEPA with several state-of-the-art SSL methods.
-For classification tasks, we use AUC (Area Under the ROC Curve) and F1 scores as our primary evaluation metrics.
-AUC evaluates model performance across all possible thresholds, providing a comprehensive and stable measure of discriminative ability.
-In contrast, the F1 score balances precision and recall at a fixed threshold.
-In our experiments, we use the AUC as our main metric, as it provides a more robust evaluation of the model’s performance across varying classification thresholds, ensuring that the results are not overly dependent on a single decision boundary.
-For multi-label classification, the AUC is computed by averaging the scores of binary classifications for each label.
-For multi-class classification, the AUC is calculated using the average of one-vs-rest approach.
-In both cases, the F1 score is computed using a macro-averaged approach that averages the F1 scores of all classes.
-In most cases, ECG-JEPA outperforms other SSL methods that leverage hand-crafted augmentations, demonstrating its effectiveness in learning general representations of ECG data.
-In experiments, E-JEPArb and E-JEPAmb refer to ECG-JEPA models trained with random masking and multi-block masking strategies, respectively.
-6.1 Linear Evaluation
-Tables 1 and 2 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets.
-We train a linear classifier on top of the frozen representations for 10 epochs, and evaluate its performance on downstream tasks.
-As shown in the tables, ECG-JEPA consistently outperforms other SSL methods, demonstrating superior efficiency and effectiveness with substantially reduced computational resources.
+In this section, we evaluate the performance of the learned representations across various downstream tasks to demonstrate their generalizability and ability to capture essential ECG features.
+ECG-JEPA is compared against several state-of-the-art self-supervised learning (SSL) methods.
+For classification tasks, we use AUC (Area Under the ROC Curve) and F1 scores as evaluation metrics.
+AUC provides a comprehensive measure of discriminative ability by considering performance across all classification thresholds, making it more robust to variations in decision boundaries.
+In contrast, the F1 score balances precision and recall at a fixed threshold, offering insights into the model’s performance when a specific decision boundary is chosen.
+In multi-label classification, we compute AUC by averaging the scores from binary classification for each label, while for multi-class classification, AUC is calculated using the one-vs-rest approach.
+For both tasks, F1 scores are macro-averaged across all classes to ensure equal weighting of each class in the final score.
+In most cases, ECG-JEPA consistently outperforms other SSL methods that rely on hand-crafted augmentations, highlighting its effectiveness in learning generalizable representations from ECG data.
+In our experiments, E-JEPArb and E-JEPAmb refer to ECG-JEPA models trained using random masking and multi-block masking strategies, respectively.
 Table 1: Linear evaluation on multi-label task.
@@
 MoCo v3111Score as reported in [22]. 800 0.739 0.142 0.712 0.080
-MTAE [^1] Score as reported in [23]. 800 0.807 0.437 0.818 0.349
-MLAE [^1] Score as reported in [23]. 800 0.779 0.382 0.794 0.263
+MTAE111 800 0.807 0.437 0.818 0.349
+MLAE111 800 0.779 0.382 0.794 0.263
 ST-MEM 800 0.888 0.566 0.973 0.805
@@
 E-JEPAmb 100 0.896 0.628 0.973 0.819
+6.1 Linear Evaluation
+Tables 1 and 2 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets.
+We train a linear classifier on top of the frozen representations for 10 epochs and evaluate its performance on downstream tasks.
+Further training beyond 10 epochs does not lead to any significant improvement in performance.
+As shown in the tables, ECG-JEPA consistently outperforms other SSL methods, demonstrating superior efficiency and effectiveness with substantially reduced computational resources.
 6.2 Fine-tuning
 Fine-tuning is another method to evaluate the quality of learned representations, as it tests the model’s ability to adapt its pre-trained features to new tasks.
-In the fine-tuning process, we add a linear classification head at the end of the encoder and train the whole network for 10 epochs.
+In the fine-tuning process, we add a linear classification head at the end of the encoder and train the entire network for 10 epochs.
+Similar to linear evaluation, training for 10 epochs is sufficient, as further training does not lead to additional performance gains.
 Fine-tuning can potentially enhance performance beyond what is achieved with linear evaluation alone.
 Table 3 presents the results of fine-tuning on the PTB-XL and CPSC2018 datasets.
-ECG-JEPA is compared with other SSL methods in a multi-class classification setting.
+ECG-JEPA is compared with other SSL methods as well as supervised methods in a multi-class classification setting, where the student network is trained directly for the supervised task.
 The results indicate that ECG-JEPA achieves the highest AUC and F1 scores on PTB-XL and the highest AUC on CPSC2018.
-These results highlight ECG-JEPA’s effectiveness in fine-tuning scenarios, demonstrating its ability to leverage transfer learning by refining and adapting its learned representations for improved performance in specific classification tasks.
+These findings underscore ECG-JEPA’s effectiveness in fine-tuning scenarios, demonstrating its ability to refine and adapt learned representations for improved performance in specific classification tasks.
 Table 3: Fine-tuning on multi-class task.
@@
 Method Epochs AUC F1 AUC F1
-MoCo v3111Score as reported in [23]. 800 0.913 0.644 0.967 0.838
-MTAE [^1] Score as reported in [23]. 800 0.910 0.613 0.961 0.769
-MLAE [^1] Score as reported in [23]. 800 0.915 0.625 0.973 0.816
-CMSC [^1] Score as reported in [23]. 800 0.877 0.510 0.938 0.717
+Supervised 100 0.887 0.608 0.893 0.566
+MoCo v3111 800 0.913 0.644 0.967 0.838
+MTAE111 800 0.910 0.613 0.961 0.769
+MLAE111 800 0.915 0.625 0.973 0.816
+CMSC111 800 0.877 0.510 0.938 0.717
 ST-MEM 800 0.929 0.668 0.977 0.820
@@
 CMSC 300 0.750 $\pm$ 0.008 0.792 $\pm$ 0.001
-CPC 100 0.523 ±$\pm 0.006$ 0.560 $\pm$ 0.005
+CPC 100 0.523 $\pm$ 0.006 0.560 $\pm$ 0.005
 E-JEPArb 100 0.836 $\pm$ 0.006 0.887 $\pm$ 0.000
@@
 6.4 Reduced Lead Evaluation
-Because transformer architectures can accept variable input lengths, we can evaluate ECG-JEPA’s performance on reduced lead tasks.
-In this experiment, we perform a linear evaluation on the PTB-XL multi-label task using reduced leads.
-Specifically, we utilize only a single lead (Lead II) and two leads (Lead II and V1), training linear classifiers on top of the learned representations for 10 epochs.
-We then compare ECG-JEPA’s performance with ST-MEM, another model supporting reduced lead evaluation.
-Table 5 presents the results of this reduced lead linear evaluation.
-Notably, ECG-JEPA performs well even in the reduced lead setting.
-This robustness is particularly valuable since most daily mobile devices typically output only one or two leads, highlighting ECG-JEPA’s potential for practical applications in mobile health monitoring.
-Table 5: Reduced lead evaluation. Linear evaluation on PTB-XL multi-label task with reduced lead settings.
+Since transformer architectures can handle variable input lengths, we evaluated ECG-JEPA’s performance with reduced leads.
+In this experiment, we conducted a linear evaluation on the PTB-XL multi-label task using only a single lead (Lead II) and two leads (Lead II and V1), training linear classifiers on the learned representations for 10 epochs [^3] We compare only with ST-MEM, as it is a transformer-based model whose pretrained weights are publicly available..
+Table 5 presents the results.
+Notably, ECG-JEPA maintains strong performance even with fewer leads, which is valuable for practical applications in mobile health monitoring, where most devices typically output only one or two leads.
+Table 5: Reduced lead evaluation. Linear evaluation of PTB-XL multi-label classification in single-leade (II) and dual-lead (II and V1) configurations.
 1-Lead 2-Lead
@@
 6.5 ECG Feature Extraction
-ECG features are essential for diagnosing and monitoring cardiac conditions.
-We evaluate the model’s ability to extract key ECG features, including heart rate and QRS duration, from the learned representations of the PTB-XL dataset.
-Unlike classification tasks, which are highly perceptual, ECG features are directly related to the morphology of the signal.
-For example, heart rate can be measured by the average distance between R-peaks, and QRS duration can be measured by the average width of the QRS complex.
-Several methods have been proposed to segment ECG signals [30, 31, 32, 33].
-Segmentation models can be used to extract ECG features.
-In this experiment, we employed a publicly available segmentation model [33] to extract ECG features from the PTB-XL dataset, which served as our ground truth labels.
-We then trained a linear regression model on the learned representations to predict these ECG features, using the mean squared error (MSE) as the loss function.
-Table 6 presents the performance comparison, showing the means and standard deviations of the absolute differences between the extracted and predicted values across all samples in the PTB-XL test set.
-Interestingly, although the model’s learned representations are implied to capture high-level features based on their strong performance in linear evaluation tasks across various datasets, they still retain the ability to recover low-level information, such as ECG features.
-This dual capability of capturing both high-level and low-level information is noteworthy and highlights the model’s versatility.
+Extracting ECG features is crucial for diagnosing and monitoring cardiac conditions.
+In this experiment, we assess the model’s ability to extract key features such as heart rate and QRS duration from the learned representations of the PTB-XL dataset.
+Unlike classification tasks, which focus on perceptual patterns, ECG features are directly tied to the signal’s morphology.
+For example, heart rate is derived from the average distance between R-peaks, while QRS duration is measured by the average width of the QRS complex.
+Various methods exist for segmenting ECG signals [28, 29, 30, 31], which can be used to extract ECG features.
+For this experiment, we utilized a publicly available segmentation model [31] to generate ground truth labels for heart rate and QRS duration from the PTB-XL dataset.
+We then trained a linear regression model on the learned representations to predict these features, using mean squared error (MSE) as the loss function.
+Table 6 shows the performance comparison, reporting the means and standard deviations of the absolute differences between the predicted and extracted values for the heart rate and QRS duration across the PTB-XL test set.
+Interestingly, although the model’s representations are designed to capture high-level features, they retain the capacity to recover low-level ECG features.
+This dual ability to encode both high-level semantics and low-level morphology underscores the versatility of ECG-JEPA, highlighting its potential in both diagnostic and real-world applications.
 Table 6: ECG feature extraction.
@@
 For a fair comparison, we trained ECG-JEPA with and without CroPA for 100 and 200 epochs and compared their performance on the PTB-XL multi-class task.
-The results show that CroPA considerably improves the model’s performance, demonstrating its effectiveness in capturing inter-lead relationships and enhancing the model’s ability to learn meaningful representations.
+The results show that CroPA improves the model’s performance, demonstrating its effectiveness in capturing inter-lead relationships and enhancing the model’s ability to learn meaningful representations.
 Table 7: Effect of CroPA. Linear evaluation (lin) and fine-tuning (ft) results on PTB-XL multi-class task with ECG-JEPArb and ECG-JEPAmb. In both cases, CroPA improves the model’s performance.
@@
 This finding is particularly significant for applications constrained by limited computational resources.
-Appendix A Supplementary Materials
-A.1 Experimental Details on ECG-JEPA
+10 Data Availability
+All datasets used in this work are publicly available, and our code can be accessed at https://github.com/sehunfromdaegu/ecg_jepa.
+11 Funding
+This work was supported by the National Research Foundation of Korea(NRF) Grant 2023R1A2C1005562, funded by the Korean government.
+Appendix A Experimental Details
+A.1 Hyperparameters for ECG-JEPA
 Hyperparameters for ECG-JEPA pretraining, linear evaluation, and fine-tuning are provided in Tables 9, 10, and 11, respectively.
@@
 epochs 10
-A.2 Pretraining Other SSL Models
+A.2 Hyperparameters for Other Pretrained Models
 Besides pretraining ECG-JEPA, we also pretrained other models, including CMSC [19], CPC [15], and SimCLR [7] using the same datasets as ECG-JEPA.
@@
 Table 12, 13, 14, and 15 show the distribution of the PTB-XL and CPSC2018 datasets in both multi-label and multi-class settings.
+Note that the sum of samples in each class exceeds the total number of ECG recordings in multi-label task.
 The PTB-XL dataset is stratified into ten folds, where the first eight folds are used for training, the ninth fold for validation, and the tenth fold for testing.
@@
 The original CPSC2018 dataset consists of 6,877 ECG recordings, but we excluded recordings with a length of less than 10 seconds, resulting in 6,867 ECG recordings.
-Table 12: PTB-XL multi-label distribution. The sum of samples in each class exceeds the total number of ECG recordings due to the multi-label nature of the dataset.
+Table 12: PTB-XL multi-label distribution.
 # ECG Norm MI STTC CD HYP
@@
 Test 2158 963 550 521 496 262
-Table 13: PTB-XL multi-class distribution. ECGs with multiple labels are excluded from the PTB-XL multi-label dataset.
+Table 13: PTB-XL multi-class distribution.
 # ECG Norm MI STTC CD HYP
@@
 Test 1650 912 256 242 184 56
-Table 14: CPSC2018 multi-label distribution. The sum of samples in each class exceeds the total number of ECG recordings due to the multi-label nature of the dataset.
+Table 14: CPSC2018 multi-label distribution.
 # ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
@@
 Table 15: CPSC2018 multi-class distribution.
-The number of ECG recordings for each superclass in the PTB-XL dataset.
-Note that the sum of samples in each class exceeds the total number of ECG recordings due to multi-label the nature of the dataset.
 # ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
@@
 Test 814 113 126 19 16 85 72 72 112 199
-A.4 Exponential Moving Average
+Appendix B Exponential Moving Average
 The teacher network is initialized as a copy of the student network and is updated using an exponential moving average (EMA) of the student’s weights. The EMA is computed as follows:
@@
 Here, $\text{ema}_{0}$ and $\text{ema}_{1}$ represent the initial and final values of the momentum parameter, respectively. For our implementation, $\text{ema}_{0}=0.996$ and $\text{ema}_{1}=1.0$.
-A.5 Sufficiency of 8 Leads in ECG Analysis
+Appendix C Sufficiency of 8 Leads in ECG Analysis
 In a standard 12-lead ECG, leads III, aVR, aVL, and aVF can be derived from leads I and II, as shown by the following equations:
@@
 This simplification not only reduces the computational load but also maintains the diagnostic integrity of the analysis.
-A.6 Comparison with 12-Lead Model
+Appendix D Comparison with 12-Lead Model
 We now investigate the practical sufficiency of using 8 leads for ECG-JEPA pretraining.

### v4

@@
 Abstract
-We propose a self-supervised learning method for 12-lead Electrocardiogram (ECG) analysis, named ECG Joint Embedding Predictive Architecture (ECG-JEPA).
-ECG-JEPA employs a masking strategy to learn semantic representations of ECG data.
-Unlike existing methods, ECG-JEPA predicts at the hidden representation level rather than reconstructing raw data.
-This approach offers several advantages in the ECG domain: (1) it avoids producing unnecessary details, such as noise, which is common in standard ECG; and (2) it addresses the limitations of naïve L2 loss between raw signals.
-Another key contribution is the introduction of a special masked attention tailored for 12-lead ECG data, Cross-Pattern Attention (CroPA).
-CroPA enables the model to effectively capture inter-patch relationships.
-Additionally, ECG-JEPA is highly scalable, allowing efficient training on large datasets.
+Electrocardiogram (ECG) captures the heart’s electrical signals, offering valuable information for diagnosing cardiac conditions.
+However, the scarcity of labeled data makes it challenging to fully leverage supervised learning in medical domain.
+Self-supervised learning (SSL) offers a promising solution, enabling models to learn from unlabeled data and uncover meaningful patterns.
+In this paper, we show that masked modeling in the latent space can be a powerful alternative to existing self-supervised methods in the ECG domain.
+We introduce ECG-JEPA, a SSL model for 12-lead ECG analysis that learns semantic representations of ECG data by predicting in the hidden latent space, bypassing the need to reconstruct raw signals.
+This approach offers several advantages in the ECG domain: (1) it avoids producing unnecessary details, such as noise, which is common in ECG; and (2) it addresses the limitations of naïve L2 loss between raw signals.
+Another key contribution is the introduction of Cross-Pattern Attention (CroPA), a specialized masked attention mechanism tailored for 12-lead ECG data.
+ECG-JEPA is trained on the union of several open ECG datasets, totaling approximately 180,000 samples, and achieves state-of-the-art performance in various downstream tasks including ECG classification and feature prediction.
 Our code is openly available at https://github.com/sehunfromdaegu/ECG_JEPA.
 1 Introduction
-Electrocardiograms (ECG) provide a non-invasive method to measure the electrical activity of the heart over time, serving as a crucial tool for diagnosing various cardiac conditions.
+Electrocardiography is a non-invasive method to measure the electrical activity of the heart over time, serving as a crucial tool for diagnosing various cardiac conditions.
 While numerous supervised methods have been developed to detect heart diseases using ECG data [1, 2, 3], these models often face significant performance degradation when applied to data distributions different from those on which they were trained.
-Self-supervised learning (SSL) has emerged as a powerful paradigm for learning general representations across various domains, including natural language processing (NLP) [4, 5, 6], computer vision (CV) [7, 8, 9], and video analysis [10, 11].
-Despite its promise, applying SSL to ECG data presents unique challenges.
+This challenge points to the need for more flexible approaches that can learn robust, transferable representations from ECG data.
+Self-supervised learning (SSL) offers an alternative approach by learning general representations in diverse domains, such as natural language processing (NLP) [4, 5, 6], computer vision (CV) [7, 8, 9], and video analysis [10, 11].
+Despite this promise, the application of SSL to ECG data presents unique challenges.
 For instance, data augmentation, which is essential in many SSL architectures, is more complex for ECG than for computer vision data.
@@
 These architectures may also miss visually subtle but diagnostically critical features, such as P-waves and T-waves, which are imperative for diagnosing certain cardiac conditions.
-In this work, we propose a novel ECG Joint-Embedding Predictive Architecture (ECG-JEPA) tailored for 12-lead ECG data, effectively addressing the aforementioned challenges.
-ECG-JEPA utilizes a transformer architecture to capture the semantic meanings of ECG.
-By masking several patches of the ECG data, ECG-JEPA predicts the abstract representations of the missing patches, indicating a high-level understanding of the data.
-Additionally, we develop a novel masked-attention for multi-lead ECG data, coined Cross-Pattern Attention (CroPA).
+In this work, we propose ECG Joint-Embedding Predictive Architecture (ECG-JEPA) tailored for 12-lead ECG data, effectively addressing the aforementioned challenges.
+ECG-JEPA utilizes a transformer architecture to capture the semantic meaning of the ECG.
+By masking several patches of the ECG, ECG-JEPA predicts abstract representations of the missing segments, indicating a high-level understanding of the data.
+Additionally, we develop a novel masked-attention for multi-lead ECG data, chich we call Cross-Pattern Attention (CroPA).
 CroPA incorporates clinical knowledge into the model as an inductive bias, guiding it to focus on clinically relevant patterns and relationships across leads.
-We demonstrate the following contributions of ECG-JEPA:
+Figure 1: 12-lead ECG with baseline wander artifact.
+Our contributions are as follows:
   • ECG-JEPA achieves notable improvements in linear evaluation and fine-tuning on classification tasks compared to existing SSL methods without hand-crafted augmentations.
@@
   • ECG-JEPA can also recover important ECG features, including heart rate and QRS duration, which are classical indicators used in ECG evaluation.
-This is the first work to demonstrate that learned representations can effectively recover various ECG features.
+This is the first work to demonstrate that learned representations can effectively recover ECG features.
   • ECG-JEPA is highly scalable, allowing efficient training on large datasets.
-For instance, ECG-JEPA is trained for only 100 epochs, yet outperforms other ECG SSL models on most downstream tasks, taking approximately 26 hours on a single RTX 3090 GPU.
-Figure 1: 12-lead ECG.
-The standard 12-lead ECG captures electrical activity from different angles around the heart, providing comprehensive cardiac assessment.
-Artifacts and noise, such as powerline interference and baseline wander, are common in ECG recordings.
-The figure shows an example of a 12-lead ECG with baseline wander.
+For instance, ECG-JEPA is trained for only 100 epochs, yet outperforms other ECG SSL models on most downstream tasks, taking approximately 22 hours on a single RTX 3090 GPU.
+In summary, ECG-JEPA introduces a robust SSL framework for 12-lead ECG analysis, overcoming traditional SSL limitations with clinically inspired design elements, scalable architecture, and demonstrated effectiveness on a wide range of tasks.
 2 Background
@@
 SSL leverages inherent data patterns to learn useful representations, allowing models to adapt to various downstream tasks with greater robustness to data imbalances [12].
-In this section, we briefly explain key SSL techniques in Sections 2.1 and 2.2, including generative architectures and joint-embedding predictive architectures.
-Additionally, we provide an overview of the ECG and its key features in Sections 2.3 and 2.4, highlighting the critical characteristics essential for understanding ECG data.
-2.1 Generative Architectures
-Generative architectures involve reconstructing an input $x$ from its degraded version $x^{\prime}$ using an encoder-decoder framework.
-The idea behind generative architectures is that the ability to reconstruct the clean data from a corrupted one indicates a good understanding of the data.
-The encoder maps the perturbed input $x^{\prime}$ into a latent representation, which the decoder then uses to generate the original input $x$ [13].
-2.2 Joint-Embedding Predictive Architectures
-Joint-Embedding Predictive Architectures (JEPA) [14] process pairs $x$ and their corrupted versions $x^{\prime}$ to obtain representations $z$ and $z^{\prime}$ through encoders.
-Unlike generative architectures that predict in the input space, JEPA predicts in the hidden representation space by reconstructing $z$ from $z^{\prime}$.
-This approach effectively avoids the challenge of predicting unpredictable details, which is common in biological signals.
+We begin in Section 2.1 with an overview of the ECG and its key features, highlighting the critical characteristics essential for understanding ECG data.
+In Sections 2.2 and 2.3, we briefly explain key SSL techniques and their specific applications to ECG, respectively.
 2.1 Electrocardiogram (ECG)
 The electrocardiogram (ECG) is a non-invasive diagnostic method that records the heart’s electrical activity over time using electrodes placed on the skin.
-The standard 12-lead ECG captures electrical activity from multiple angles.
+The standard 12-lead ECG captures electrical activity of the heart from multiple angles.
 These 12 leads are categorized into limb leads (I, II, III), augmented limb leads (aVR, aVL, aVF), and chest leads (V1-V6).
@@
 Refer to Figure 1 for an illustration.
-In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads (see C).
+[…]
+In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads following the Einthoven’s law [13]:
+[…]
+$\text{aVR}=-(\text{I}+\text{II})/2$
+$\text{aVL}=(\text{I}-\text{II})/2$
+$\text{aVF}=(\text{II}-\text{I})/2.$
 This choice maintains the necessary diagnostic information while optimizing computational efficiency.
 Figure 2: Key ECG Features.
-2.4 ECG Features
-[…]
-Figure 3: ECG-JEPA.
-The teacher network processes the complete ECG data, generating fully contextualized representations.
-In contrast, the student network receives only the non-masked patches.
-The student’s representations are then passed to the predictor, which reconstructs the fully contextualized representations produced by the teacher.
-The predicted 6 representations are then compared with the corresponding teacher’s representation.
-This figure illustrates $L=3$ leads with $N=5$ subintervals and $Q=3$ unmasked subintervals.
+Figure 3: ECG-JEPA training overview. For illustration, we use $L=3$, $N=5$ subintervals and $Q=3$ unmasked subintervals.
+2.2 Self-Supervised Learning Architectures
+Self-supervised learning can be broadly categorized into contrastive and non-contrastive methods.
+Non-contrastive methods can be further divided into generative and non-generative architectures.
+For a broader introduction to SSL, see [14].
+In contrastive learning, the model is encouraged to produce similar representations for semantically related inputs $x^{{}^{\prime}}$ and $x^{{}^{\prime\prime}}$, while pushing apart the representations of unrelated inputs $x^{{}^{\prime}}$ and $y^{{}^{\prime}}$.
+SimCLR [7] is one of the most popular contrastive methods, using two different augmentations of a single input $x$ to form semantically similar pairs $x^{{}^{\prime}}$ and $x^{{}^{\prime\prime}}$.
+Beyond contrastive methods, generative architectures have been particularly successful in recent large language models [4, 5, 6] and in computer vision [8].
+Generative architectures involve reconstructing a sample $x$ from its degraded version $x^{\prime}$ using an encoder-decoder framework.
+The premise is that reconstructing clean data from a corrupted version reflects the model’s deep understanding of the underlying data structure.
+The encoder maps the perturbed input $x^{\prime}$ into a latent representation, which the decoder then uses to reconstruct the original input $x$ [15].
+Recently, the authors of [16] observed that generative architectures prioritize learning principal subspaces of the data, which may limit their capacity to capture semantic representations for perceptual tasks.
+As an alternative, non-generative methods have shown promise across domains, including computer vision [17, 18, 19, 9] and video analysis [11].
+Among these, the Joint-Embedding Predictive Architecture (JEPA) [20] processes an input pair $x$ and its corrupted versions $x^{\prime}$ to obtain representations $z$ and $z^{\prime}$ through encoders.
+Unlike generative architectures that make predictions in the input space, JEPA performs prediction in the latent space by reconstructing $z$ from $z^{\prime}$.
+This approach effectively avoids the challenge of predicting unpredictable details, a common issue in biological signals.
+2.3 Related Works
+Several studies have worked on capturing semantically meaningful representations of 12-lead ECG data.
+Contrastive Multi-segment Coding (CMSC) [21] splits an ECG into two segments, encouraging similar representations for compatible segments while separating incompatible ones.
+Contrastive Predictive Coding (CPC) [22], applied in [23], predicts future ECG representations in a contrastive manner, but its reliance on LSTM modules makes it inefficient for large datasets.
+More recently, [24] introduced masked autoencoders for ECG, proposing temporal and channel masking strategies, Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE).
+Similarly, [25] proposed ST-MEM, which masks random time intervals for each lead.
+However, both MLAE and ST-MEM may struggle with the high correlations between ECG leads, potentially oversimplifying the prediction task.
 3 Methodology
@@
 Compared to the standard self-attention mechanism, which treat all patches equally, CroPA reflects a structured approach that mirrors the process of multi-lead signal interpretation, leading to improved performance in downstream tasks.
-4 Related Work
-This section provides an overview of existing SSL methods that have been applied to ECG data. We compare ECG-JEPA with these methods on various downstream tasks in Section 6.
-4.1 Contrastive Learning
-Contrastive learning has been widely explored in various research works for its effectiveness in unsupervised representation learning.
-Contrastive Predictive Coding (CPC) [15] aims to capture useful data representations by predicting future observations in the latent space.
-It uses sequential data where the “true future" is treated as the positive sample, and other observations within the same sequence are treated as negative samples.
-SimCLR [7] and MoCo [16] also leverage contrastive approaches, with positive samples being augmented versions of the same image and negative samples being other images.
-SimCLR relies on large batch sizes and extensive data augmentation, while MoCo uses a momentum encoder to maintain a dynamic dictionary of encoded samples, enabling more stable training with smaller batch sizes.
-4.2 Generative Architectures
-Generative architectures, first popularized by BERT [4] in natural language processing (NLP), have since inspired a variety of approaches in other domains.
-For example, in computer vision, Masked Autoencoders (MAE) [8] extend this concept by masking portions of an input image and training the model to reconstruct the missing parts, similar to how BERT predicts masked tokens in text.
-Recently, the authors of [17] pointed out that generative architectures often prioritize learning the principal subspaces of the data, which may limit their effectiveness in capturing semantic representations for perceptual tasks.
-Figure 4: Cross-Pattern Attention (CroPA). The patch in the middle attends only to the colored patches, which are in the same lead and temporal space.
-4.3 Joint-Embedding Predictive Architectures
-In contrast to generative approaches, which focus on reconstructing data in the input space, joint-embedding predictive architectures (JEPA) aim to predict in the representation space, allowing models to learn more abstract and semantic features.
-An example of this approach is Data2vec [18], which introduced an SSL architecture that predicts the average of the encoder’s final layer outputs.
-Other examples include I-JEPA [9] and V-JEPA [11], which predict the hidden representations of masked patches, enabling semantic representation learning for images and videos, respectively.
-4.4 Self-Supervised Learning for ECG
-Various SSL models have been explored for 12-lead ECG data.
-Contrastive Multi-segment Coding (CMSC) [19] splits an ECG into two segments, encouraging similar representations for compatible segments while separating incompatible ones.
-Contrastive Predictive Coding (CPC) [15], applied in [20], predicts future ECG representations, but its reliance on LSTM modules makes it inefficient for large datasets.
-More recently, [21] introduced masked autoencoders for ECG, proposing temporal and channel masking strategies, Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE).
-Similarly, [22] proposed ST-MEM, which masks random time intervals for each lead.
-However, both MLAE and ST-MEM may struggle with the high correlations between ECG leads, making some predictions too easy.
+Figure 4: Cross-Pattern Attention (CroPA). The patch in the middle attends only to the colored patches.
+Figure 5: Squares following the encoder represent the representations of ECG patches.
+The representations are subsequently averaged through a pooling layer, with the resulting vector (highlighted in cyan) serving as an abstract representation of the ECG data.
+3.4 ECG representation
+After training, we use only the student network as the encoder.
+The encoder outputs are average-pooled to obtain the final ECG representation, which serves as the feature vector for downstream tasks.
+See Figure 5 for an illustration.
 4 Experimental Settings
@@
 5 Experiments
+Table 1: Linear evaluation on multi-label and multi-class tasks. Our proposed method outperforms all baselines, achieving the highest AUC and F1 scores across both tasks and datasets.
+Multi-label Task Multi-class Task
+Method Epochs PTB-XL CPSC2018 PTB-XL CPSC2018
+AUC F1 AUC F1 AUC F1 AUC F1
+ST-MEM 800 0.896 0.662 0.964 0.752 0.888 0.566 0.973 0.805
+SimCLR 300 0.866 0.624 0.890 0.523 0.842 0.496 0.918 0.624
+CMSC 300 0.802 0.472 0.767 0.206 0.796 0.442 0.787 0.391
+CPC 100 0.620 0.167 0.687 0.091 0.600 0.201 0.672 0.210
+MoCo v31 800 - - - - 0.739 0.142 0.712 0.080
+MTAE1 800 - - - - 0.807 0.437 0.818 0.349
+MLAE1 800 - - - - 0.779 0.382 0.794 0.263
+ECG-JEPArb 100 0.906 0.690 0.969 0.769 0.894 0.616 0.974 0.805
+ECG-JEPAmb 100 0.912 0.712 0.971 0.789 0.896 0.628 0.973 0.819
+  • 1
+    Scores reported in [25]; results for multi-label tasks were not available.
 In this section, we evaluate the performance of the learned representations across various downstream tasks to demonstrate their generalizability and ability to capture essential ECG features.
@@
 For both tasks, F1 scores are macro-averaged across all classes to ensure equal weighting of each class in the final score.
-In most cases, ECG-JEPA consistently outperforms other SSL methods that rely on hand-crafted augmentations, highlighting its effectiveness in learning generalizable representations from ECG data.
-In our experiments, E-JEPArb and E-JEPAmb refer to ECG-JEPA models trained using random masking and multi-block masking strategies, respectively.
-Table 1: Linear evaluation on multi-label task.
-PTB-XL CPSC2018
-Method Epochs AUC F1 AUC F1
-ST-MEM 800 0.896 0.662 0.964 0.752
-SimCLR 300 0.866 0.624 0.890 0.523
-CMSC 300 0.802 0.472 0.767 0.206
-CPC 100 0.620 0.167 0.687 0.091
-E-JEPArb 100 0.906 0.690 0.969 0.769
-E-JEPAmb 100 0.912 0.712 0.971 0.789
-Table 2: Linear evaluation on multi-class task.
-PTB-XL CPSC2018
-Method Epochs AUC F1 AUC F1
-MoCo v3111Score as reported in [22]. 800 0.739 0.142 0.712 0.080
-MTAE111 800 0.807 0.437 0.818 0.349
-MLAE111 800 0.779 0.382 0.794 0.263
-ST-MEM 800 0.888 0.566 0.973 0.805
-SimCLR 300 0.842 0.496 0.918 0.624
-CMSC 300 0.796 0.442 0.787 0.391
-CPC 100 0.600 0.201 0.672 0.210
-E-JEPArb 100 0.894 0.616 0.974 0.805
-E-JEPAmb 100 0.896 0.628 0.973 0.819
+In most cases, ECG-JEPA consistently outperforms other SSL methods that rely on hand-crafted augmentations, highlighting its effectiveness in learning generalizable representations.
+In our experiments, ECG-JEPArb and ECG-JEPAmb refer to ECG-JEPA models trained using random masking and multi-block masking strategies, respectively.
 5.1 Linear Evaluation
-Tables 1 and 2 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets.
+Table 1 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets.
 We train a linear classifier on top of the frozen representations for 10 epochs and evaluate its performance on downstream tasks.
@@
 Fine-tuning is another method to evaluate the quality of learned representations, as it tests the model’s ability to adapt its pre-trained features to new tasks.
-In the fine-tuning process, we add a linear classification head at the end of the encoder and train the entire network for 10 epochs.
+We add a linear classification head at the end of the encoder and train the entire network for 10 epochs.
 Similar to linear evaluation, training for 10 epochs is sufficient, as further training does not lead to additional performance gains.
@@
 Table 2 presents the results of fine-tuning on the PTB-XL and CPSC2018 datasets.
-ECG-JEPA is compared with other SSL methods as well as supervised methods in a multi-class classification setting, where the student network is trained directly for the supervised task.
+ECG-JEPA is compared with other SSL methods as well as supervised methods in a multi-class classification setting, where the student network is trained directly from the scratch.
 The results indicate that ECG-JEPA achieves the highest AUC and F1 scores on PTB-XL and the highest AUC on CPSC2018.
-These findings underscore ECG-JEPA’s effectiveness in fine-tuning scenarios, demonstrating its ability to refine and adapt learned representations for improved performance in specific classification tasks.
-Table 3: Fine-tuning on multi-class task.
-PTB-XL CPSC2018
-Method Epochs AUC F1 AUC F1
-Supervised 100 0.887 0.608 0.893 0.566
-MoCo v3111 800 0.913 0.644 0.967 0.838
-MTAE111 800 0.910 0.613 0.961 0.769
-MLAE111 800 0.915 0.625 0.973 0.816
-CMSC111 800 0.877 0.510 0.938 0.717
-ST-MEM 800 0.929 0.668 0.977 0.820
-SimCLR 300 0.905 0.650 0.934 0.693
-CPC [^2] We did not fine-tune CPC due to its slow training process. 100 - - - -
-E-JEPArb 100 0.944 0.710 0.980 0.821
-E-JEPAmb 100 0.937 0.680 0.983 0.799
 5.3 Low-shot Linear Evaluation
@@
 This suggests that ECG-JEPA can be particularly well-suited for transfer learning where labeled data is scarce.
-Table 4: Low-shot linear evaluation on the multi-label PTB-XL task.
-The mean and standard deviation of macro AUC are reported for 1% (192 samples) and 10% (1923 samples) of the training set, selected three times independently.
+Table 2: Fine-tuning on multi-class task.
+Method Epochs PTB-XL CPSC2018
+AUC F1 AUC F1
+Supervised 100 0.887 0.608 0.893 0.566
+MoCo v31 800 0.913 0.644 0.967 0.838
+MTAE1 800 0.910 0.613 0.961 0.769
+MLAE1 800 0.915 0.625 0.973 0.816
+CMSC1 800 0.877 0.510 0.938 0.717
+ST-MEM 800 0.929 0.668 0.977 0.820
+SimCLR 300 0.905 0.650 0.934 0.693
+CPC2 100 - - - -
+ECG-JEPArb 100 0.944 0.710 0.980 0.821
+ECG-JEPAmb 100 0.937 0.680 0.983 0.799
+  • 1
+    Scores reported in [25].
+  • 2
+    We did not fine-tune CPC due to its slow training process.
+Table 3: Low-shot linear evaluation on the multi-label PTB-XL.
+The mean and standard deviation of macro AUCs are reported for 1% (192 samples) and 10% (1923 samples) of the training set, selected three times independently.
 PTB-XL
@@
 CPC 100 0.523 $\pm$ 0.006 0.560 $\pm$ 0.005
-E-JEPArb 100 0.836 $\pm$ 0.006 0.887 $\pm$ 0.000
-E-JEPAmb 100 0.843 $\pm$ 0.004 0.894 $\pm$ 0.003
+ECG-JEPArb 100 0.836 $\pm$ 0.006 0.887 $\pm$ 0.000
+ECG-JEPAmb 100 0.843 $\pm$ 0.004 0.894 $\pm$ 0.003
 5.4 Reduced Lead Evaluation
@@
 Notably, ECG-JEPA maintains strong performance even with fewer leads, which is valuable for practical applications in mobile health monitoring, where most devices typically output only one or two leads.
-Table 5: Reduced lead evaluation. Linear evaluation of PTB-XL multi-label classification in single-leade (II) and dual-lead (II and V1) configurations.
+Table 4: Reduced lead evaluation. Linear evaluation of PTB-XL multi-label classification in single-leade (II) and dual-lead (II and V1).
 1-Lead 2-Lead
@@
 ST-MEM 0.832 0.571 0.848 0.597
-E-JEPArb 0.846 0.596 0.877 0.647
-E-JEPAmb 0.849 0.593 0.880 0.657
+ECG-JEPArb 0.846 0.596 0.877 0.647
+ECG-JEPAmb 0.849 0.593 0.880 0.657
 5.5 ECG Feature Extraction
@@
 Unlike classification tasks, which focus on perceptual patterns, ECG features are directly tied to the signal’s morphology.
-For example, heart rate is derived from the average distance between R-peaks, while QRS duration is measured by the average width of the QRS complex.
 Various methods exist for segmenting ECG signals [31, 32, 33, 34], which can be used to extract ECG features.
@@
 This dual ability to encode both high-level semantics and low-level morphology underscores the versatility of ECG-JEPA, highlighting its potential in both diagnostic and real-world applications.
-Table 6: ECG feature extraction.
-Means and standard deviations of absolute differences between the predicted and extracted values for heart rate and QRS duration.
-Small values indicate better performance.
-The final row provides the overall mean and standard deviation for both features across all samples in the PTB-XL test set.
+Table 5: ECG feature prediction results on PTB-XL multi-lable test set. The mean heart rate and QRS duration in the test set are 70.01 BPM (±$\pm 17.65$) and 90.48 ms (±$\pm 17.02$), respectively.
+Mean Absolute Error
 Method Heart Rate (BPM) QRS Dur. (ms)
@@
 CPC 11.40 $\pm$ 11.04 11.55 $\pm$ 11.55
-E-JEPArb 1.54 $\pm$ 2.62 4.81 $\pm$ 4.29
-E-JEPAmb 1.45 $\pm$ 2.44 4.41 $\pm$ 4.08
-Mean 75.01 $\pm$ 17.65 90.48 $\pm$ 17.02
+ECG-JEPArb 1.54 $\pm$ 2.62 4.81 $\pm$ 4.29
+ECG-JEPAmb 1.45 $\pm$ 2.44 4.41 $\pm$ 4.08
 6 Ablation Study
@@
 The results show that CroPA improves the model’s performance, demonstrating its effectiveness in capturing inter-lead relationships and enhancing the model’s ability to learn meaningful representations.
-Table 7: Effect of CroPA. Linear evaluation (lin) and fine-tuning (ft) results on PTB-XL multi-class task with ECG-JEPArb and ECG-JEPAmb. In both cases, CroPA improves the model’s performance.
+Table 6: Effect of CroPA. Linear evaluation (lin) and fine-tuning (ft) results on PTB-XL multi-class task.
 lin ft
@@
 Multi-block (0.175, 0.225) 4 0.912 0.712
-8 Future Work
-Our method has potential applications in various physiological multivariate signals, such as EEG and EMG.
-These signals share characteristics with ECG, including their multivariate nature and high dimensionality, making them suitable for our proposed approach.
-Another promising direction involves the integration of multi-modal data.
-For example, multi-lead ECG data could be combined with other diagnostic modalities, such as chest X-rays (CXR), to provide a more comprehensive understanding of a patient’s condition.
-One significant challenge in pursuing these extensions is the scarcity of large-scale datasets.
-Addressing this limitation is crucial for advancing and validating our method across diverse applications.
+Table 8: Comparison of 8-Lead and 12-Lead Models on PTB-XL multi-label.
+Model epochs AUC F1
+8-Lead 100 0.906 0.690
+12-Lead 100 0.905 0.699
+6.3 Comparison with 12-Lead Model
+We now investigate the practical sufficiency of using 8 leads for ECG-JEPA pretraining.
+To evaluate the impact of this reduction, we trained models using both 8 leads and 12 leads and compared their performance on the linear evaluation of a multi-label task for PTB-XL.
+Table 8 presents the results of this comparison using ECG-JEPArb.
+As expected, the performance difference between the 8-lead and 12-lead models is minimal, indicating that using 8 leads is sufficient for effective pretraining without significant loss of information.
+7 Discussion
+7.1 Insights and Interpretations
+The results demonstrate that ECG-JEPA effectively captures high-quality representations from 12-lead ECG signals, as evidenced by its superior performance across various downstream tasks, including classification, low-shot learning, and feature extraction.
+The model’s ability to maintain robust performance under reduced lead configurations underscores its practical applicability in resource-constrained scenarios, such as wearable devices and remote health monitoring.
+Moreover, the proposed Cross-Pattern Attention (CroPA) mechanism introduces a clinically inspired inductive bias, aligning with the physiological patterns of multi-lead ECG signals.
+This targeted attention contributes to enhanced model performance, particularly in tasks requiring inter-lead correlations.
+The findings validate the importance of incorporating domain-specific design elements into self-supervised learning frameworks for medical data.
+Compared to previous SSL approaches, ECG-JEPA offers significant advancements.
+While several methods rely on extensive augmentations or manual feature engineering, ECG-JEPA bypasses these requirements by learning semantic representations directly in the latent space.
+To the best of our knowledge, we are the first to demonstrate that ECG representations learned through self-supervised learning can successfully recover key ECG features such as heart rate and QRS duration.
+This finding highlights the dual capability of ECG-JEPA to encode both high-level semantic information and low-level morphological details, making it versatile for various diagnostic and monitoring tasks.
+These results pave the way for further exploration of self-supervised learning methods in uncovering clinically meaningful patterns in physiological signals.
+7.2 Limitations and Challenges
+While ECG-JEPA achieves state-of-the-art performance, certain limitations remain.
+One notable limitation is the lack of inherent explainability in the model’s learned representations.
+Although ECG-JEPA effectively captures semantic and morphological features, it provides limited insights into how these features are utilized for specific predictions, which can be crucial in medical applications.
+The absence of a clear interpretability mechanism may hinder its adoption in clinical settings, where understanding the decision-making process is often as important as the results themselves.
+7.3 Broader Implications
+The implications of this work extend beyond ECG analysis.
+The principles underlying ECG-JEPA, particularly the combination of latent-space prediction and domain-specific attention mechanisms, could inspire advancements in other multivariate physiological signal domains, such as EEG and EMG.
+By leveraging these principles, researchers could develop models capable of extracting meaningful representations from diverse biomedical data, potentially accelerating progress in multimodal diagnostic systems.
+7.4 Future Directions
+Looking ahead, integrating ECG-JEPA with complementary diagnostic modalities, such as chest X-rays or echocardiograms, could provide a more holistic understanding of cardiac health.
+This multi-modal approach has the potential to improve diagnostic accuracy by leveraging the strengths of different data types, enabling a richer representation of patient conditions.
+One significant challenge in pursuing these extensions is the scarcity of large-scale datasets in other modalities.
+Addressing this limitation is crucial for advancing the multi-model foundation model.
 8 Conclusion
@@
 This finding is particularly significant for applications constrained by limited computational resources.
-10 Data Availability
-All datasets used in this work are publicly available, and our code can be accessed at https://github.com/sehunfromdaegu/ecg_jepa.
-11 Funding
-This work was supported by the National Research Foundation of Korea(NRF) Grant 2023R1A2C1005562, funded by the Korean government.
 Appendix A Experimental Details
-A.1 Hyperparameters for ECG-JEPA
-Hyperparameters for ECG-JEPA pretraining, linear evaluation, and fine-tuning are provided in Tables 9, 10, and 11, respectively.
+A.1 Downstream Datasets Details
+Table 9, and 10 show the distribution of the PTB-XL and CPSC2018 datasets, respectively. Note that the sum of samples in each class exceeds the total number of ECG recordings in multi-label task.
+The PTB-XL dataset is stratified into ten folds, where the first eight folds are used for training, the ninth fold for validation, and the tenth fold for testing.
+In our experiments, we used the first nine folds for training and the tenth fold for testing, as we did not observe overfitting during linear evaluation and fine-tuning.
+For the CPSC2018 dataset, only the training set is publicly available, which is stratified into seven folds.
+We used the first six folds for training and the seventh fold for testing, omitting the validation set.
+The original CPSC2018 dataset consists of 6,877 ECG recordings, but we excluded recordings with a length of less than 10 seconds, resulting in 6,867 ECG recordings.
+Table 9: PTB-XL Distribution.
+Type Set # ECG Norm MI STTC CD HYP
+Multi-label Total 21799 9514 5469 5235 4898 2649
+Train 19230 8551 4919 4714 4402 2387
+Test 2158 963 550 521 496 262
+Multi-class Total 16244 9069 2532 2400 1708 535
+Train 14594 8157 2276 2158 1524 479
+Test 1650 912 256 242 184 56
+Table 10: CPSC2018 Distribution.
+Type Set # ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
+Multi-label Total 6867 918 1220 235 220 721 614 699 868 1854
+Train 5989 805 1059 206 197 632 534 615 742 1616
+Test 878 113 161 29 23 89 80 84 126 238
+Multi-class Total 6391 918 975 178 185 685 531 606 783 1530
+Train 5577 805 849 159 169 600 459 534 671 1331
+Test 814 113 126 19 16 85 72 72 112 199
+A.2 Hyperparameters for ECG-JEPA
+Hyperparameters for ECG-JEPA pretraining, linear evaluation, and fine-tuning are provided in Tables 11, 12, and 13, respectively.
 In ECG-JEPAmb, the number of visible patches in ECG-JEPAmb varies more than in ECG-JEPArb, resulting in higher GPU memory usage.
@@
 epochs 10
-A.2 Hyperparameters for Other Pretrained Models
+A.3 Hyperparameters for Other Pretrained Models
 Besides pretraining ECG-JEPA, we also pretrained other models, including CMSC [21], CPC [22], and SimCLR [7] using the same datasets as ECG-JEPA.
@@
 These augmentations aimed to enhance the robustness of the model to various signal distortions.
-A.3 Downstream Datasets Details
-Table 12, 13, 14, and 15 show the distribution of the PTB-XL and CPSC2018 datasets in both multi-label and multi-class settings.
-Note that the sum of samples in each class exceeds the total number of ECG recordings in multi-label task.
-The PTB-XL dataset is stratified into ten folds, where the first eight folds are used for training, the ninth fold for validation, and the tenth fold for testing.
-In our experiments, we used the first nine folds for training and the tenth fold for testing, as we did not observe overfitting during linear evaluation and fine-tuning.
-For the CPSC2018 dataset, only the training set is publicly available, which is stratified into seven folds.
-We used the first six folds for training and the seventh fold for testing, omitting the validation set.
-The original CPSC2018 dataset consists of 6,877 ECG recordings, but we excluded recordings with a length of less than 10 seconds, resulting in 6,867 ECG recordings.
-Table 12: PTB-XL multi-label distribution.
-# ECG Norm MI STTC CD HYP
-Multi-label 21799 9514 5469 5235 4898 2649
-Train 19230 8551 4919 4714 4402 2387
-Test 2158 963 550 521 496 262
-Table 13: PTB-XL multi-class distribution.
-# ECG Norm MI STTC CD HYP
-Multi-class 16244 9069 2532 2400 1708 535
-Train 14594 8157 2276 2158 1524 479
-Test 1650 912 256 242 184 56
-Table 14: CPSC2018 multi-label distribution.
-# ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
-Multi-label 6867 918 1220 235 220 721 614 699 868 1854
-Train 5989 805 1059 206 197 632 534 615 742 1616
-Test 878 113 161 29 23 89 80 84 126 238
-Table 15: CPSC2018 multi-class distribution.
-# ECG Norm PVC AF LBBB STE 1AVB PAC STD RBBB
-Multi-class 6391 918 975 178 185 685 531 606 783 1530
-Train 5577 805 849 159 169 600 459 534 671 1331
-Test 814 113 126 19 16 85 72 72 112 199
 Appendix B Exponential Moving Average
@@
 Here, $\text{ema}_{0}$ and $\text{ema}_{1}$ represent the initial and final values of the momentum parameter, respectively. For our implementation, $\text{ema}_{0}=0.996$ and $\text{ema}_{1}=1.0$.
-Appendix C Sufficiency of 8 Leads in ECG Analysis
-In a standard 12-lead ECG, leads III, aVR, aVL, and aVF can be derived from leads I and II, as shown by the following equations:
-[…]
-$\text{aVR}=-\frac{(\text{I}+\text{II})}{2}$
-$\text{aVL}=\text{I}-\frac{\text{II}}{2}$
-$\text{aVF}=\text{II}-\frac{\text{I}}{2}$
-These derivations demonstrate that leads III, aVR, aVL, and aVF do not provide additional independent information beyond what is captured by leads I and II.
-Therefore, using only 8 leads (I, II, V1, V2, V3, V4, V5, and V6) can be sufficient for effective ECG analysis.
-This simplification not only reduces the computational load but also maintains the diagnostic integrity of the analysis.
-Appendix D Comparison with 12-Lead Model
-We now investigate the practical sufficiency of using 8 leads for ECG-JEPA pretraining.
-To evaluate the impact of this reduction, we trained models using both 8 leads and 12 leads and compared their performance on the linear evaluation of a multi-label task for PTB-XL.
-Table 16 presents the results of this comparison using ECG-JEPArb.
-As expected, the performance difference between the 8-lead and 12-lead models is minimal, indicating that using 8 leads is sufficient for effective pretraining without significant loss of information.
-Table 16: Comparison of 8-Lead and 12-Lead Models.
-Model epochs AUC F1
-8-Lead 100 0.906 0.690
-12-Lead 100 0.905 0.699
+Acknowledgment
+The author would like to thank Otto van Koert for his introduction to electrocardiography and for providing valuable insights on the subject.